A steel I-beam used in the construction of a bridge is 9.2m long at 20°C. On a hot summer day, the temperature rises to 40°C. What is the change in the beam's length due to linear expansion?


1
Expert's answer
2021-05-12T17:55:44-0400

"\Delta L=\alpha L \Delta T,"

"\Delta L=12\cdot 10^{-6}\cdot 9.2\cdot 20=2.2~mm."